the highest point of a roller coaster is 29.6 meters. a cart with its occupants has a mass of 23000kg and a gravitational potential energy of 667,865.0 Jewels at this location. the cart races down to the bottom of a hill (point b) which is almost at ground level. what is the velocity of this cart at point b


1
Expert's answer
2021-03-03T09:37:39-0500
"KE_a+PE_a=KE_b+PE_b,""0+PE_a=KE_b+0,""PE_a=\dfrac{1}{2}mv_b^2,""v_b=\sqrt{\dfrac{2PE_a}{m}}=\sqrt{\dfrac{2\cdot6678650\ J}{23000\ kg}}=24.1\ \dfrac{m}{s}."